OpenStudy (twizttiez):

Variable y varies directly with x^2, and y = 72 when x = 6. Enter the constant of variation for the quadratic variation equation. y = ____ x^2

OpenStudy (twizttiez):

@MrNood

OpenStudy (twizttiez):

I think my answer is 2

OpenStudy (mrnood):

you are looking for an equation y=kx^2 you know that y = 72 when x = 6 so write y=kx^2 using 72 for y, and 6 for x

OpenStudy (twizttiez):

y varies directly with x^2 y= k x^2 , where k is any constant y=72; x=6 72 = k(6)^2 72 = 36 k k = 2 The constant of variation is 2 Right?

OpenStudy (mrnood):

correct

OpenStudy (twizttiez):

Ok i have a question that i really 100% don't know what to do can you help?

OpenStudy (twizttiez):

Do you mind if i put it here or would you prefer i open a new ?